OpenStudy (anonymous):

Write an equation of a line in slop intercept form that is parallel to y= 3x + 6 and passes through (10,2.5)

OpenStudy (irishboy123):

if it's parallel, then y=mx+c becomes y=3x+c, ie slope = 3. if it pass through (10, 2.5), then 2.5 = 3(10) + c c = 2.5-30 = 27.5 y = 3x + 27.5

OpenStudy (anonymous):

Thank you so much!!